OpenStudy (anonymous):

Evaluate 3x − 5y when x = 3 and y = −2

OpenStudy (anonymous):

That would be: 3*(3) - 5*(-2) = 9 + 10 = 19

OpenStudy (anonymous):

positive right?

OpenStudy (anonymous):

Exactly what Saikam did! And yes positive

OpenStudy (anonymous):

ok